A homeowner estimates that it will take him 13 days to roof his house. A professional roofer estimates that he could roof the house in 6 days. How long will it take if the homeowner helps the roofer?

Answer by checkley79(3341)

1/6+1/13=1/X
(13+6)/6*13=1/X
19/78=1/X19X=78
X=78/19
X=4.1 DAYS WORKING TOGETHER.
PROOF:
1/6+1/13=1/4.1
(13+6)/6*13=1/4.1
19/78=1/4.1
78=19*4.1
78=78